# Evaluating Privacy Perceptions, Experience, and Behavior of Software Development Teams

### v1

@@ Abstract
With the increase in the number of privacy regulations, small development teams are forced to make privacy decisions on their own. In this paper, we conduct a mixed-method survey study, including statistical and qualitative analysis, to evaluate the privacy perceptions, practices, and knowledge of members involved in various phases of software development (SDLC). Our survey includes 362 participants from 23 countries, encompassing roles such as product managers, developers, and testers. Our results show diverse definitions of privacy across SDLC roles, emphasizing the need for a holistic privacy approach throughout SDLC. We find that software teams, regardless of their region, are less familiar with privacy concepts (such as anonymization), relying on self-teaching and forums. Most participants are more familiar with GDPR and HIPAA than other regulations, with multi-jurisdictional compliance being their primary concern. Our results advocate the need for role-dependent solutions to address the privacy challenges, and we highlight research directions and educational takeaways to help improve privacy-aware software development.
@@ 1 Introduction
The number of privacy violations in the US rose from $\sim$447 in 2012 to $\sim$2100 in 2023 [73], with $\sim$45of companies experiencing harmful breaches [69]. In response, many countries adopted new privacy regulations [46] such as the European General Data Protection Regulation (GDPR) [18]. With these new regulations, developers are under increased scrutiny to implement privacy engineering solutions throughout the Software Development Life Cycle (SDLC) or face financial penalties. Many mobile apps are initially developed by a small team of independent developers with limited privacy expertise or access to legal/policy resources to make privacy decisions [5, 6, 43]. Without mandatory privacy training and a low barrier to entering the market, developers may lack essential knowledge of privacy concepts, leading them to overlook their privacy responsibilities prior to their apps’ deployment.
@@
In recent years, several frameworks and approaches, including Privacy by Design (PbD), are introduced to help developers incorporate privacy rules throughout the SDLC [21, 23, 22, 76, 45, 57, 11, 29, 30, 12, 37, 35]. However, few works examined the implementation of these solutions from the developers’ perspective and their impact on privacy practices. Tahaei et al. [65] suggest that the lack of mainstream adoption of PbD causes confusion for software teams. Recent studies show that developers mainly see privacy as a burden and afterthought and are not familiar with basic privacy concepts [17, 58, 59, 6, 13, 66, 25, 63, 43, 8, 26]. These studies focus on only a limited group of developers and overlook the broader SDLC roles and the unique challenges faced by each specific role. They do not also examine how factors such as legal expertise, privacy regulations, and regional differences influence software teams’ privacy perceptions and practices.
@@
In this paper, we conduct a large mixed-method survey study on Prolific with 189 participants located in the US and 173 participants located in 22 other countries (in total 362), who are involved in various roles in the SDLC – including administrators (e.g., scrum masters, product managers), development and Quality Assurance (QA) teams, and information security/privacy experts. The non-US participants are located in EU+UK (132), South Africa (21), Canada/Mexico (18), and South America (2). Our goal is to identify the current state of privacy comprehension, obstacles, practices, and behaviors in various SDLC roles, and the privacy gaps that have yet to be addressed. Our survey comprises three parts: pre-screening questions (e.g., describing their product/customers), generic questions regarding participants’ demographics (education, role, company size, etc.); and role-specific questions to examine their perceptions, experiences, and behaviors. We combine the participants’ demographics (e.g., location data) with the role-specific responses to help determine:
@@
– RQ1: Are there any differences in privacy perceptions among various roles and demographics (e.g., education, company size, or location)?
@@
– RQ2: Does access to privacy experts (e.g., a Chief Privacy Officer (CPO)) impact privacy perceptions and practices?
@@
Our results show that participants have diverse perceptions/definitions of privacy, showcasing the need for a refined approach to privacy in SDLC. Scrum masters, product managers, and information security/privacy experts define privacy more in terms of limited disclosure, while developers and QAs perceive privacy as control over personal information. Our study finds a lack of adoption of most PbD strategies and other privacy techniques, such as Privacy Enhancing Technologies (PETs) and Privacy Impact Assessment (PIA) in SDLC. Most QA members rely on legal/privacy experts to ensure the protection of users’ data, and they lack privacy knowledge and expertise. Members of software teams are generally self-taught regarding privacy concepts, and most are not familiar with regulations that exist in the US, such as the California Consumer Privacy Act (CCPA) [24] and the Children’s Online Privacy Protection Rule (COPPA) [19]. We also find that software teams face challenges in both understanding and adhering to privacy regulations, especially across multiple jurisdictions. These findings highlight the need for more privacy-focused education and training. Comparing regional-specific trends regarding the use of PETs, the creation of PIAs, or the presence of a CPO, we did not observe any differences among our participants regardless of their location. This shows that privacy practices are primarily determined by the culture of the organization, and are not influenced by various regulations across regions [4]. Our results highlight the challenges faced in various SDLC roles and advocate the need for role-dependent solutions to address them. Based on these findings, we outline research directions and educational takeaways to help improve privacy-aware software development.
@@ 2 Related Work
Understanding developers’ privacy expertise and concerns has been explored in research through user studies with developers and analysis of developers’ forums [42, 34, 13, 68, 32]. Tahaei et al. [62] conducted 12 semi-structured interviews with developers to examine their strategies to protect users’ privacy. They identified factors such as poor privacy culture, tensions between privacy and other business rules, and lack of standardized privacy tools impact how developers implement privacy. They also emphasize the role of privacy champions to minimize such barriers. Iwaya et al. [34] interviewed 30 practitioners from nine countries, uncovering three key themes: personal privacy mindset, organizational privacy aspects, and privacy engineering practices. They highlight the regulations’ impact on practitioners’ behaviors and corporate cultures, and the need for a privacy-conscious environment. Alomar and Egelman [3] focused on how children’s app developers handle compliance. Through surveys and interviews with 127 participants, they identify that developers often rely on app markets to spot privacy issues and they struggle with third-party tools to maintain compliance. Balebako et al. [6] examined how mobile app developers make privacy and security decisions. Through interviews with 13 developers and a survey of 228 US-based developers, they reveal that smaller companies tend to exhibit fewer positive privacy and security behaviors. Their research emphasizes the need for simplified, cost-effective privacy tools such as privacy checklists, especially for small firms. This study was done in 2014 when regulations such as GDPR and CCPA did not exist. Khandelwal et al. [39] demonstrate that implementing and maintaining privacy labels pose challenges that require resources, tools, and guidelines.
@@
Other works analyze developers’ forums to understand their challenges. The analysis of Stack Overflow (SO) posts shows that developers frequently query regarding PbD, compliance, and confidentiality  [68, 61]. Delile et al. [15] compared privacy questions on SO with responses generated by ChatGPT to identify whether ChatGPT could be used as an alternative tool. Their results indicate in some cases ($\sim$30%), SO is more accurate than ChatGPT responses. Li et al. [41] and Parsons et al. [49] studied posts on several Reddit forums and identified most discussions on personal data usage occur in response to external events such as Android OS changes or privacy laws.
@@
Although our work complements these efforts, it is the first study to assess privacy perceptions, practices, and knowledge of members of software teams involved in various roles in SDLC through a large-scale mixed-method approach. Prior work focused only on developers (i.e., programmers) in the US and a few countries, whereas we study members from various SDLC (including product managers, QA, etc.) roles spanning 23 countries. In our work, we investigate how factors such as organizational aspects (e.g., the presence of a CPO) and participants’ demographics (e.g., role, education, and location) impact privacy perceptions, experience, and behaviors of software teams. We also explore how frequently developers use online forums for privacy-related queries.
@@ 3 Study Design
In this paper, we aim to understand how members of software teams in small, medium-sized, and large companies (i.e., with <20, 21–100, and 100+ employees), implement privacy in their software applications and examine their level of privacy comprehension, expertise, practices and behaviors based on various demographics (such as roles, location, education level, etc.). For this purpose, we first conducted a pilot study to evaluate our survey design and then a large-scale study with members of software teams in 23 countries. Our pilot study was completed in January 2023, while our large-scale study was done between February–April 2023.
@@ 3.1 Survey Tools
We utilized Qualtrics for survey creation, a platform supported by our university. Using Qualtrics, we customized our survey to individualize questions based on the participants’ role (Q9) [^1] Survey questions: http://tinyurl.com/2p9n49e4 as defined in Table 1. For example, we asked developers about familiarity with PbD (Q39) and their use of forums such as Reddit (Q32), while information security members were asked about the management of access control, encryption algorithms, and certificates (Q60-Q62).
@@
Survey Platform Selection We conducted the large-scale survey using Qualtrics integration on the Prolific [48, 67] platform since it provides a higher pay rate and allows selecting from a more specific pool of participants with basic programming knowledge, in our case - software teams.
Tahaei et al. and Kaur et al. [67, 38] recommend using Prolific and MTurk for large-scale surveys. Although pre-screening via programming questions is recommended [67, 38, 14, 53], it has limitations: (a) overusing such questions could lead to automatically responding correctly without paying attention to the questions [67]; (b) in studies such as ours where the software teams include a variety of roles (e.g., product manager, QA, etc.) as well as with specific programming skills (e.g., JavaScript developers), having programming knowledge questions may bias the participants’ pools towards more experienced developers in larger companies with traditional programming knowledge, preventing recruiting novice developers and those in other SDLC roles; and (c) AI tools like ChatGPT [10] are widely accessible and can handle code-based questions, thus, these questions are no longer a strong barrier to screen participants. Our analysis of Danilova et al.’s [14] pre-screening questions with GPT-3.5 shows that the tool can answer the questions with 95% accuracy. We discuss how we mitigate these issues below and in Section 4.
Conducting the Survey In our survey, participants were required to be at least 18 years old, fluent in English, and working in industries such as Graphic Design, Information Services and Data Processing, Product Development, Software, Video Games, etc. We paid an average of \$25.17/hr to participants who completed the survey. We filtered participants based on their industry and verified their eligibility on Prolific. Out of the 686 participants who started the survey, 14 did not give their consent, and 295 did not finish the survey; hence, were excluded from our analysis. We then conducted another filtering process, to ensure that the participants work in the software industry and have in fact software development experience. We asked them “In short, tell us about your product and who are your customers.” and manually evaluated their responses. We found that the majority of them are involved in software development activities such as “I make a productivity app for mac & windows to record & share the user’s screen”. We eliminated 15 participants as we could not verify their involvement in SDLC; for example, those with responses as “NA” or “I sell home decor items. My customers are primarily women.” Following these steps, we ended up with a total of 362 participants for our final count.
@@ 3.2 Pilot Survey: University Students
Our pilot survey participants were our university’s graduate students over the age of 18 from the disciplines of Computing and Information Science, Electrical and Computer Engineering, and Business, who had experience in software development, IT, or related fields. To maintain their anonymity, we did not collect any personally identifiable information such as their contact, names, or company names.
The goal of the pilot study was to gather initial insights and feedback before the deployment of our main study on Prolific. Upon the IRB approval, we launched the survey using Qualtrics. The survey consisted of 40 questions, including 13 short and 27 multiple-choice questions. We estimated that the survey takes $\sim$30-40 minutes to complete. Every participant was presented with the same set of questions regardless of their role on a software team. We used the responses to improve our large-scale survey (i.e., Subsection 3.3).
@@ 3.3 Software Teams Survey
The main feedback we received from the pilot study was that the survey required too much time to complete ($\sim$27 minutes). To address this limitation and to focus on capturing participants’ perspectives related to their SDLC roles, we separated the survey questions according to the roles. This shortened the survey duration by 12 minutes and enhanced the quality of the responses we received. We first asked all participants the same set of 10 questions that are partly related to demographics and the degree of privacy understanding (e.g., how do they define privacy). We then divided the remainder of the questions into four groups, one for each role defined in Table 1. Our breakdown loosely follows the SDLC phases, but we separated the ISec roles from the SD roles to evaluate the significance of security or privacy knowledge in our survey.
@@ 3.4 Survey Questions
The survey includes a mix of demographic, perception, experiential, and behavioral questions to answer our research questions (see Section 1). The complete list of questions (except questions 1-3, which are the required Prolific identification questions and our consent form) is found here. [^2] Survey questions: http://tinyurl.com/2p9n49e4
Demographic Questions These questions collect basic information about the participants such as age, education, their SDLC role, and the company size. e.g.,“What areas/roles of the development team are you currently involved with?”.
Perception Questions These questions aim to understand participants’ perceptions toward privacy; e.g., “How do you define privacy?”.
Experiential Questions These questions ask about the participants’ experience with privacy-related tasks and challenges; e.g., “What was the process for the Privacy Impact Assessment, and who was involved?”.
Behavioral Questions These questions ask about the participants’ behaviors and knowledge related to privacy; e.g., “List any privacy-by-design strategies you have used or know:”.

### v2

@@ Abstract
With the increase in the number of privacy regulations, small development teams are forced to make privacy decisions on their own. In this paper, we conduct a mixed-method survey study, including statistical and qualitative analysis, to evaluate the privacy perceptions, practices, and knowledge of members involved in various phases of the Software Development Life Cycle (SDLC). Our survey includes 362 participants from 23 countries, encompassing roles such as product managers, developers, and testers. Our results show diverse definitions of privacy across SDLC roles, emphasizing the need for a holistic privacy approach throughout SDLC. We find that software teams, regardless of their region, are less familiar with privacy concepts (such as anonymization), relying on self-teaching and forums. Most participants are more familiar with GDPR and HIPAA than other regulations, with multi-jurisdictional compliance being their primary concern. Our results advocate the need for role-dependent solutions to address the privacy challenges, and we highlight research directions and educational takeaways to help improve privacy-aware SDLC.
Copyright is held by the author/owner. Permission to make digital or hard copies of all or part of this work for personal or classroom use is granted without fee.
USENIX Symposium on Usable Privacy and Security (SOUPS) 2024.
August 11–13, 2024, Philadelphia, PA, United States.
@@ 1 Introduction
With the vast increase in privacy violations in the US and around the world [95], many countries have adopted new privacy regulations [66], such as the European General Data Protection Regulation (GDPR) [27]. With these new regulations, developers are under increased scrutiny while implementing privacy engineering solutions throughout the Software Development Life Cycle (SDLC) or face financial penalties. Many mobile apps are initially developed by a small team of independent developers with limited privacy expertise or access to legal/policy resources to make privacy decisions [6, 7, 63]. Research shows that this lack of access to privacy expertise leads to challenges in creating concise, accurate and consistent privacy policies [99, 77, 78, 65, 11, 13, 96, 72, 52], implementing privacy concepts throughout the SDLC - from early analysis to testing [26, 40, 86, 63], and distinguishing between privacy and security approaches, tools and regulations [81, 82, 7, 21, 89, 36, 9, 40].
@@
In recent years, several approaches, including Privacy by Design (PbD), have been introduced to help developers incorporate privacy rules throughout the SDLC [31, 33, 32, 99, 65, 80, 17, 44, 45, 20, 55, 53]. However, few works examined the implementation of these solutions from the developers’ perspective and their impact on privacy practices. Most studies focus on only a limited group of developers and overlook the broader SDLC roles and the unique challenges faced by each role (e.g., product manager when defining privacy requirements or the QAs when identifying privacy leaks) [61, 51, 21, 91, 48]. They also do not examine how factors such as legal expertise, regulations, and regional differences influence software teams’ privacy perceptions and practices.
@@
In this paper, we conduct a large mixed-method survey study on Prolific with 189 participants located in the US and 173 participants located in 22 other countries (in total 362), who are involved in various roles in the SDLC – including administrators (e.g., scrum masters, product managers), development and Quality Assurance (QA) teams, and information security/privacy experts. The non-US participants are located in EU+UK (132), South Africa (21), Mexico (15), Canada (3), and South America (2). Our goal is to identify the current state of privacy comprehension, practices, and behaviors in various SDLC roles, and the privacy gaps that have yet to be addressed. Our survey comprises of three parts: pre-screening questions (e.g., describing their product/customers), generic questions regarding participants’ demographics (education, role, company size, etc.); and role-specific questions to examine their perceptions, experiences, and behaviors. We combine the participants’ demographics (e.g., location data) with the role-specific responses to help determine:
@@
– RQ1: Are there any differences in privacy perceptions among various roles, locations, and other demographics?
@@
– RQ2: Does access to privacy experts (e.g., a Chief Privacy Officer - CPO) impact privacy perceptions and practices?
@@
Our results show that participants have diverse perceptions/definitions of privacy, showcasing the need for a refined approach to privacy in SDLC. Scrum masters, product managers, and information security/privacy experts define privacy more in terms of limited disclosure, while developers and QAs perceive privacy as control over personal information. Our study finds a lack of adoption of most PbD strategies and other privacy techniques, such as Privacy Enhancing Technologies (PETs) and Privacy Impact Assessment (PIA), in SDLC. Most QA members rely on legal/privacy experts to protect users’ data, and they lack privacy knowledge and expertise. Members of software teams are generally self-taught regarding privacy concepts, and most are not familiar with regulations that exist in the US, such as the California Consumer Privacy Act (CCPA) [35] and the Children’s Online Privacy Protection Rule (COPPA) [29]. We also find that software teams face challenges in both understanding and adhering to privacy regulations, especially across multiple jurisdictions. These findings highlight the need for more privacy-focused education and training. Comparing regional-specific trends regarding the use of PETs, the creation of PIAs, or the presence of a CPO, we did not observe any differences among our participants, regardless of their location. This shows that privacy practices are primarily determined by the culture of the organization and are not influenced by various regulations across regions [5]. Our results highlight the challenges faced in various SDLC roles and advocate the need for role-dependent solutions to address them. Based on these findings, we outline research directions and educational takeaways to help improve privacy-aware SDLC.
@@ 2 Related Work
Understanding developers’ privacy expertise and concerns has been explored in research through user studies with developers and analysis of developers’ forums [61, 51, 21, 91, 48]. Tahaei et al. [85] and Horstmann et al. [48] conducted interviews with developers and privacy experts and identified factors such as poor privacy culture, tensions between privacy and other business rules, lack of proper communication between privacy experts and developers, lack of standardized privacy tools, and mismatch between the technical expertise of developers and privacy experts that impact how developers implement privacy. They also emphasize the role of privacy champions to minimize such barriers. In 2014 (i.e., pre-GDPR and CCPA), Balebako et al. [7] examined how app developers make privacy and security decisions and revealed that smaller companies exhibit fewer positive privacy and security behaviors. Their research emphasizes the need for simplified, cost-effective privacy tools such as privacy checklists, especially for small firms. Other studies [51, 3, 57, 48] with practitioners and developers highlight that while regulations impact practitioners’ behaviors and corporate cultures, the developers and practitioners mostly rely on app markets to spot privacy issues, and they struggle with implementing and maintaining privacy labels, as well as leveraging third-party tools to maintain compliance.
@@
The analysis of Stack Overflow (SO) posts shows that developers frequently query regarding PbD, compliance, and confidentiality  [91, 84]. Delile et al. [24] compared privacy questions on SO with responses generated by ChatGPT to identify whether ChatGPT could be used as an alternative tool. Their results indicate that, in $\sim$30of cases SO is more accurate than ChatGPT. Li et al. [59] and Parsons et al. [70] studied posts on several Reddit forums and identified that most discussions on personal data usage occur in response to external events such as Android OS changes or privacy laws.
@@
These studies pinpoint developers’ challenges in correctly implementing privacy requirements and maintaining compliance. Our work complements these efforts; however, it is the first study to assess privacy perceptions, practices, and knowledge of members of software teams involved in various roles in SDLC through a large-scale mixed-method approach. Prior work focused only on developers (i.e., programmers) in the US and a few countries, whereas we studied members from various SDLC roles (including product managers, QA, etc.) spanning 23 countries. In our work, we investigate how factors such as organizational aspects (e.g., the presence of a CPO) and participants’ demographics (e.g., role, education, and location) impact privacy perceptions, experience, and behaviors of software teams. We also explore how frequently developers use online forums for privacy-related queries.
@@ 3 Study Design
In this paper, we aim to understand how members of software teams in small, medium-sized, and large companies (i.e., with <20, 21–100, and 100+ employees), implement privacy in their software applications and examine their level of privacy comprehension, expertise, practices, and behaviors based on various demographics (such as roles, location, education level, etc.). For this purpose, we first conducted a pilot study to evaluate our survey design and then a large-scale study with members of software teams in 23 countries. Our pilot study was completed in January 2023, while our large-scale study was done between February–April 2023.
@@ 3.1 Survey Tools
We utilized Qualtrics for survey creation, a platform supported by our university. Using Qualtrics, we customized our survey to individualize questions based on the participants’ role (Q9) as defined in Table 1. For example, we asked developers about familiarity with PbD (Q39) and their use of forums such as Reddit (Q32), while information security members were asked about the management of access control, encryption algorithms, and certificates (Q60-Q62).
@@
Survey Platform Selection We conducted the large-scale survey using Qualtrics integration on the Prolific [68, 90] platform, since it provides a higher pay rate and allows selecting from a more specific pool of participants with basic programming knowledge, in our case - software teams.
Tahaei et al. and Kaur et al. [90, 56] recommend using Prolific and MTurk for large-scale surveys. Although pre-screening via programming questions is recommended [90, 56, 23, 74], it has limitations: (a) overusing such questions could lead to automatically responding correctly without paying attention to the questions [90]; (b) in studies such as ours where the software teams include a variety of roles (e.g., product manager, QA, etc.) as well as with specific programming skills (e.g., JavaScript developers), having programming knowledge questions may bias the participants’ pools towards more experienced developers in larger companies with traditional programming knowledge, preventing recruiting novice developers and those in other SDLC roles; and (c) AI tools like ChatGPT [15] are widely accessible and can handle code-based questions. Thus, these questions are no longer a strong barrier to screen participants. Our analysis of Danilova et al.’s [23] pre-screening questions with GPT-3.5 shows that the tool can answer the questions with 95% accuracy. We discuss how we mitigate these issues below and in Section 4.
Conducting the Survey Prolific maintains a pool of active participants who are regularly screened and vetted by the platform. In our survey, we decided on the sample size based on Prolific’s guidelines (a minimum of 300 for a representative sample). We initially pre-screened the Prolific participants based on the following requirements: (a) to be at least 18 years old, (b) fluent in English, and (c) working in industries such as Graphic Design, Information Services, Data Processing, Product Development, Software, Video Games, etc. We used their industry (rather than their role) as a filter, since Prolific does not allow selecting participants based on role. We paid an average of \$25.17/hr to those who completed the survey. After the initial pre-screening, we recruited 686 participants across both US and non-US pools. Out of the 686 participants who started the survey, 14 did not give their consent, and 295 did not finish the survey; hence, they were excluded from our analysis. We then conducted another filtering process to ensure that the participants work in the software industry and, in fact, have software development experience. We asked them, “Q4. In short, tell us about your product and who your customers are.” We manually evaluated their responses and cross-checked them with Q6 (their post-secondary degree) and Q9 (their roles). We found that most of them are involved in software development activities such as “I make a productivity app for Mac & Windows to record & share the user’s screen.” We eliminated 15 participants as we could not verify their involvement in SDLC; for example, those with responses as “NA” or “I sell home decor items. My customers are primarily women.” Following these steps, we ended up with a total of 362 participants for our final count.
@@ 3.2 Pilot Survey: University Students
Our pilot survey participants were our university’s graduate students (who mostly have industry experience through internships and part-/full-time jobs) over the age of 18 from the disciplines of Computing and Information Science, Electrical and Computer Engineering, and Business, who had experience in software development, IT, or related fields. To maintain their anonymity, we did not collect any personally identifiable information such as their contact, names, or company names.
The goal of the pilot study was to gather initial insights and feedback before the deployment of our main study on Prolific. Upon the IRB approval, we launched the survey using Qualtrics. The survey consisted of 40 questions, including 13 short and 27 multiple-choice questions, which were derived based on our informal discussions with developers in small companies and prior gaps in research. We estimated that the survey takes $\sim$30-40 minutes to complete. Every participant was presented with the same set of questions regardless of their role on a software team. We used the responses to improve our large-scale survey (i.e., Subsection 3.3).
@@ 3.3 Software Teams Survey
The main feedback we received from the pilot study was that the survey required too much time to complete ($\sim$27 minutes). To address this limitation and to focus on capturing participants’ perspectives related to their SDLC roles, we separated the survey questions according to the roles. This shortened the survey duration by 12 minutes and enhanced the quality of the responses we received. We first asked all participants the same set of 10 questions that are partly related to demographics and the degree of privacy understanding. We then divided the remainder of the questions into four groups, one for each role defined in Table 1. Our breakdown loosely follows the SDLC phases, but we separated the Information Security/Privacy (ISec) roles from the Software Developer (SD) roles to evaluate the significance of security or privacy knowledge in our survey. Although “Others” role was an option, none of the participants selected it.
@@ 3.4 Survey Questions
The survey includes a mix of demographic, perception, experiential, and behavioral questions which are crafted based on our RQs (see Section 1) and the challenges identified in prior research regarding creating privacy-preserving applications, such as understanding privacy concepts [40, 9, 2], knowledge of regulations and establishing compliance [4, 28, 33], creating consistent and accurate privacy policies [99, 12, 60, 77, 34, 69, 71, 72, 52, 96], knowledge of privacy approaches and existing tools [38, 39, 16, 40]. The complete list of questions (except questions 1-3, which are the required Prolific identification questions and our consent form) is found here. [^1] Survey questions: http://tinyurl.com/2p9n49e4
Demographic questions collect basic information about the participants, such as age, education, their SDLC role, and the company size; e.g., “What areas/roles of the development team are you currently involved with?”.
Perception questions aim to understand participants’ perceptions toward privacy; e.g., “How do you define privacy?”.
Experiential questions ask about their experience with privacy challenges and tasks; e.g., “What was the process for the Privacy Impact Assessment, and who was involved?”.
Behavioral questions ask about the participants’ behaviors and knowledge related to privacy; e.g., “List any privacy-by-design strategies you have used or know.